Convert each of the following as indicated.
1. 0.003 to percent
2. 0.035% to decimal
3. 3/5 to percent
4. 35% to reduced to fraction
5. 3.11 to percent
6. 29.5% to decimal
7. 5/4 to percent
8. 28% to a reduced fraction
Solve
1. Find 22% of 350
2. 5% of what is 240?
3. Find what percent of 40 is 12.5.
4. Find 135% of 400.
5. 120% of what is 350?
6. Find what percent of 25 is 40.
Thank You very much for your time we, you are very much appreciated.

Answer by lotanez(23)

1. if 0.3=30% (30/100)
and 0.03=3% (3/100)
then 0.003=0.03% (.03/100)
2. 0.035/100= 3.5
3. 3/5 = 0.6 (just use your caluculator and type 3 divided by 5)
4. 35/100 = 7/20
5. 3.11 x 100 = 311%
6. 29.5/100= 0.295
7. 5/4=1.25 (100)= 125%
8. 28/100= 28/4= 7; 100/4= 25 therefore, 28/100= 7/25
Solve
1. 22% of 350
350/22= 15.91, then 350-15.91=334.09
2. 5% of 240
240/5=48, 240-48=192
3. 12.5/40 = 3.2%
4. 135/100= 1.35%
5. 120/350= 2.92
6. 25/40= 1.6%
Hope this helps.